Kim and Sandy are each knitting a scarf. Kim can knit 3 rows in 5 minutes, while Sandy can knit 4 rows in 6 minutes. They start knitting at the same time and do not take any breaks. Kim wants her scarf to be 84 rows long. Sandy wants her scarf to be 88 rows long. Who will finish her scarf first and by how many minutes?
Answer by stanbon(57347)

Kim and Sandy are each knitting a scarf.
Kim can knit 3 rows in 5 minutes,
Kim rate: 3/5 row/min
----------------------------
while Sandy can knit 4 rows in 6 minutes.
Sandy rate: 2/3 row/min
----------------------------
They start knitting at the same time and do not take any breaks.
Kim wants her scarf to be 84 rows long.
Kim time for 84 rows: 84/(3/5) = 5*28 = 140 minutes
-------------------------------------------------------
Sandy wants her scarf to be 88 rows long. Who will finish her scarf first and by how many minutes?
Sandy time for 88 rows: 88/(2/3) = 3*44 = 132 minutes
=================
Cheers,
Stan H.